# DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition

### v5

@@ 2 Related Works
NER is the task of identifying named entities in text and classifying them into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of training data to achieve strong performance, which limits their application in low-resource scenarios. Some researchers apply LLMs to address this issue. GPT-NER Wang et al. (2023) employs the GPT-3 model and re-frames the task as single-entity labeling, supporting few-shot/zero-shot learning. It achieves comparable performance to supervised methods in traditional scenarios and excels in low-resource scenarios. PromptNER Ashok and Lipton (2023) achieves state-of-the-art (SOTA) accuracy in datasets with complex classification Liu et al. (2021); Ding et al. (2021) with GPT-4 and Chain-of-Thought (CoT) Wei et al. (2022), yet performs significantly worse than GPT-NER and supervised methods in classical NER datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). Furthermore, several studies apply LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024), focusing on science and medicine. Their performances surpass supervised methods in those domains, further highlighting the potential of LLMs in low-resource and complex NER tasks.
@@ 4.2 Extraction
After conducting extensive experiments, we find that the extractor’s precision consistently exceeds recall, regardless of the model or dataset, indicating that while correct entities are effectively identified, there is a tendency for under-detection. To mitigate this issue, we introduce a union strategy in result fusion Ganaie et al. (2022), allowing multiple extraction for one sentence and taking the union of the results to maximize recall. For cases of entity nesting, where different extraction rounds produce overlapping or nested entities, we apply a length-first strategy, retaining the longer entity, as longer entities generally carry more semantic meaning Nguyen and Cao (2010). For example, "Boston University" is semantically more accurate than "Boston" in the context of "She studies in Boston University". The formula of our strategy is shown below:
@@ 5 Experiment
G-1.5B 47.6 39.7 38.0 37.6 37.3 41.2 35.7 36.1 36.9 32.2 31.9 30.5 30.3 35.8 31.9 32.6
G-7B 52.3 46.4 44.8 44.8 45.7 48.1 42.3 42.1 42.7 37.3 38.2 36.8 36.5 41.1 37.3 38.6
G-GPT 60.6 57.3 56.5 55.6 55.9 58.4 54.9 53.8 49.2 46.9 47.5 47.2 47.0 48.9 47.7 48.3
P-1.5B 23.2 20.8 18.5 16.3 17.5 22.7 18.0 17.3 20.5 17.9 16.2 15.9 16.1 19.9 16.0 15.9
P-7B 44.3 35.8 33.2 32.5 31.9 40.4 37.4 35.6 39.8 33.2 32.1 31.8 31.5 37.8 35.6 34.5
P-GPT 53.0 50.5 51.2 47.9 50.2 52.3 48.7 48.5 49.4 48.5 47.1 46.6 46.0 47.4 44.1 44.0
C-1.5B 62.8 55.7 52.8 51.1 48.8 58.9 54.1 52.7 49.7 44.1 44.0 43.4 42.9 48.5 43.1 43.8
C-7B 68.2 61.5 55.3 52.9 51.4 64.5 58.8 55.3 55.7 49.9 49.7 46.5 46.1 52.9 50.2 50.0
C-GPT 73.1 67.1 67.8 66.9 67.6 68.3 67.4 67.9 61.3 57.4 56.9 56.2 56.0 59.7 56.8 56.4
@@ 5.2 DynamicNER Experiment
G-1.5B 50.2 43.5 40.4 39.8 39.3 44.1 38.9 38.7
G-7B 55.1 48.2 47.2 44.0 48.1 50.9 44.8 44.5
G-GPT 62.4 58.3 57.9 56.8 56.9 60.4 57.3 55.9
P-1.5B 21.6 18.6 17.1 14.9 15.8 20.7 16.4 15.9
P-7B 41.1 32.9 31.0 30.7 30.3 47.4 35.6 29.6
P-GPT 49.7 47.7 48.2 45.9 46.6 48.6 45.7 45.4
C-1.5B 67.6 59.9 57.9 55.7 53.5 64.0 58.5 55.1
C-7B 73.8 65.5 60.3 59.6 61.4 69.8 65.3 62.7
C-GPT 77.1 71.7 69.9 70.3 70.8 74.3 72.4 70.9
@@ Appendix D Detail of DynamicNER
Although DynamicNER’s data volume is significantly smaller than existing multilingual NER datasets, it aligns with our primary goal: evaluating LLM-based NER in fine-grained, low-resource, and dynamic scenarios. In such contexts, data volume may not be decisive. We acknowledge that smaller data volumes may lead to higher variance and overfitting. To show the robustness of our work, we conduct additional experiments on existing datasets with the model finetuned on DynamicNER. As shown in Appendix E, especially Table 7 & 8, CascadeNER, trained on DynamicNER, demonstrates strong few-shot performance on other established benchmarks. This indicates that the models have learned robust, generalizable features, avoiding overfitting.
@@ E.2 Baselines
For supervised methods, we adopt ACE+document-context by Wang et al. (2020) (SOTA of CoNLL2003) and BERT-MRC+DSC by Li et al. (2019b) (SOTA of Ontonotes 5.0 Pradhan et al. (2013)) for English datasets, while XLM-RoBERTa Conneau et al. (2020) and GEMNET by Meng et al. (2021); Fetahu et al. (2022) (SOTA of MultiCoNER) for multilingual datasets. For LLM-based methods, we adopt GPT-NER and PromptNER with GPT-4o.
@@ E.4 Experimental Results
As shown in Table 7 and 8, the results indicate that in low-resource scenarios, LLM-based methods achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music and FewNERD, and PAN-X and MultiCoNER in some languages, achieving new SOTA performance and highlighting its exceptional generalization and capability to handle complex entity categorization. However, when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X, indicating that supervised methods are still useful in some scenarios.

### v6

@@ 2 Related Works
NER is the task of identifying named entities in text and classifying them into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019), currently dominate this task. They generally rely on large amounts of training data to achieve strong performance, which limits their application in low-resource scenarios. Some researchers apply LLMs to address this issue. GPT-NER Wang et al. (2023) employs the GPT-3 model and re-frames the task as single-entity labeling, supporting few-shot/zero-shot learning. It achieves comparable performance to supervised methods in traditional scenarios and excels in low-resource scenarios. PromptNER Ashok and Lipton (2023) achieves state-of-the-art (SOTA) accuracy in datasets with complex classification Liu et al. (2021); Ding et al. (2021) with GPT-4 and Chain-of-Thought (CoT) Wei et al. (2022), yet performs significantly worse than GPT-NER and supervised methods in classical NER datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). Furthermore, several studies apply LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024), focusing on science and medicine. Their performances surpass supervised methods in those domains, further highlighting the potential of LLMs in low-resource and complex NER tasks.
@@ 4.2 Extraction
After conducting extensive experiments, we find that the extractor’s precision consistently exceeds recall, regardless of the model or dataset, indicating that while correct entities are effectively identified, there is a tendency for under-detection. To mitigate this issue, we introduce a union strategy in result fusion Ganaie et al. (2022), allowing multiple extraction for one sentence and taking the union of the results to maximize recall. For cases of entity nesting, where different extraction rounds produce overlapping or nested entities, we apply a length-first strategy, retaining the longer entity, as longer entities generally carry more semantic meaning Finkel and Manning (2009). For example, "Boston University" is semantically more accurate than "Boston" in the context of "She studies in Boston University". The formula of our strategy is shown below:
@@ 5 Experiment
GPT-NER-1.5B 47.6 39.7 38.0 37.6 37.3 41.2 35.7 36.1 36.9 32.2 31.9 30.5 30.3 35.8 31.9 32.6
GPT-NER-7B 52.3 46.4 44.8 44.8 45.7 48.1 42.3 42.1 42.7 37.3 38.2 36.8 36.5 41.1 37.3 38.6
GPT-NER-GPT 60.6 57.3 56.5 55.6 55.9 58.4 54.9 53.8 49.2 46.9 47.5 47.2 47.0 48.9 47.7 48.3
PromptNER-1.5B 23.2 20.8 18.5 16.3 17.5 22.7 18.0 17.3 20.5 17.9 16.2 15.9 16.1 19.9 16.0 15.9
PromptNER-7B 44.3 35.8 33.2 32.5 31.9 40.4 37.4 35.6 39.8 33.2 32.1 31.8 31.5 37.8 35.6 34.5
PromptNER-GPT 53.0 50.5 51.2 47.9 50.2 52.3 48.7 48.5 49.4 48.5 47.1 46.6 46.0 47.4 44.1 44.0
CascadeNER-1.5B 62.8 55.7 52.8 51.1 48.8 58.9 54.1 52.7 49.7 44.1 44.0 43.4 42.9 48.5 43.1 43.8
CascadeNER-7B 68.2 61.5 55.3 52.9 51.4 64.5 58.8 55.3 55.7 49.9 49.7 46.5 46.1 52.9 50.2 50.0
CascadeNER-GPT 73.1 67.1 67.8 66.9 67.6 68.3 67.4 67.9 61.3 57.4 56.9 56.2 56.0 59.7 56.8 56.4
@@ 5.2 DynamicNER Experiment
GPT-NER-1.5B 50.2 43.5 40.4 39.8 39.3 44.1 38.9 38.7
GPT-NER-7B 55.1 48.2 47.2 44.0 48.1 50.9 44.8 44.5
GPT-NER-GPT 62.4 58.3 57.9 56.8 56.9 60.4 57.3 55.9
PromptNER-1.5B 21.6 18.6 17.1 14.9 15.8 20.7 16.4 15.9
PromptNER-7B 41.1 32.9 31.0 30.7 30.3 47.4 35.6 29.6
PromptNER-GPT 49.7 47.7 48.2 45.9 46.6 48.6 45.7 45.4
CascadeNER-1.5B 67.6 59.9 57.9 55.7 53.5 64.0 58.5 55.1
CascadeNER-7B 73.8 65.5 60.3 59.6 61.4 69.8 65.3 62.7
CascadeNER-GPT 77.1 71.7 69.9 70.3 70.8 74.3 72.4 70.9
@@ Appendix D Detail of DynamicNER
Although DynamicNER’s data volume is significantly smaller than existing multilingual NER datasets, it aligns with our primary goal: evaluating LLM-based NER in fine-grained, low-resource, and dynamic scenarios. In such contexts, data volume may not be decisive. We acknowledge that smaller data volumes may lead to higher variance and overfitting. To show the robustness of our work, we conduct additional experiments on existing datasets with the model finetuned on DynamicNER. As shown in Appendix E, especially Table 8 & 9, CascadeNER, trained on DynamicNER, demonstrates strong few-shot performance on other established benchmarks. This indicates that the models have learned robust, generalizable features, avoiding overfitting.
@@ E.2 Baselines
For supervised methods, we adopt ACE+document-context by Wang et al. (2020) (SOTA of CoNLL2003) and BERT-MRC+DSC by Li et al. (2020) (SOTA of Ontonotes 5.0 Pradhan et al. (2013)) for English datasets, while XLM-RoBERTa Conneau et al. (2020) and GEMNET by Meng et al. (2021); Fetahu et al. (2022) (SOTA of MultiCoNER) for multilingual datasets. For LLM-based methods, we adopt GPT-NER and PromptNER with GPT-4o.
@@ E.3 Dataset
XLM-RoBERTa 92.3 59.0 65.9 72.1 70.8 66.9 80.5 64.1
ACE+doc-context 94.6 17.2 22.6 23.8 35.1 32.3 83.3 70.4
BERT-MRC+DSC 93.5 63.2 67.8 74.5 76.1 68.7 86.7 74.1
PromptNER 84.2 64.8 74.44 84.2 78.6 72.6 76.5 35.6
GPT-NER 73.5 58.0 61.2 60.8 62.4 55.8 70.0 58.4
CascadeNER (zero) 88.2 68.9 71.7 79.3 80.5 73.6 73.4 67.0
CascadeNER (few) 92.8 75.8 75.2 83.2 82.4 77.1 84.5 75.9
@@ E.4 Experimental Results
As shown in Table 8 and 9, the results indicate that in low-resource scenarios, LLM-based methods achieve significantly better results. CascadeNER surpasses existing methods on CrossNER except Music and FewNERD, and PAN-X and MultiCoNER in some languages, achieving new SOTA performance and highlighting its exceptional generalization and capability to handle complex entity categorization. However, when handling NER tasks with ample training resources and simple classifications, LLM-based methods still lag behind existing methods, whether on the English-only CoNLL2003 or the multilingual PAN-X, indicating that supervised methods are still useful in some scenarios.